OpenStudy (anonymous):

slope = -2, (5,4) is on the line

OpenStudy (anonymous):

wut iz da quiston

OpenStudy (anonymous):

find the slope intercept form

OpenStudy (shaik0124):

y-y1=m(x-x1) y-4=-2(x-5) y-4=-2x+10

OpenStudy (shaik0124):

solve further and move all terms to right hand side except y u will get slope intercept form

OpenStudy (anonymous):

thanks!!!!

OpenStudy (shaik0124):

y-4+4=-2x+10+4 y=-2x+14 is answer

OpenStudy (anonymous):

k so u do dees